In Nyeri town there are only two milk processors. The local inverse demand for milk is given by: Q = 120− P, where P denotes price, Q denotes the total quantity measured in cartons. Both milk processors have the same cost function given by C = 30Q, where C is total cost and Q is output measured in cartons. Calculate the profit for each firm







1
Expert's answer
2021-12-09T09:42:34-0500

Solution

Demand function

"Q=120-P"

So,

"P=120-Q" (1)

Since there are two firms in the market:

"Q=q_1+q_2"


Cost Function

"C=30Q"

Firm 1 Cost: "30q_1" (2)

Firm 2 Cost: "30q_1" (3)


Revenue: "Q*P"

Firm 1 Revenue:

"R= q_1 (120-(q_1+q_2))=120q_1-q_1^2-q_1q_2" (4)

Firm 2 Revenue:

"R= q_2 (120-(q_1+q_2))=120q_2-q_2q_1-q_2^2" (5)


Profit: "Revenue -Total Cost=R-C"

Firm 1 Profit: (4) minus (2)

"\u03c0(q_1,q_2)=(120q_1-q_1^2-q_1q_2)-30q_1"

"\u03c0(q_1,q_2)=90q_1-q_1^2-q_1q_2" (6)

Firm 2 profit:

"\u03c0(q_1,q_2)=(120q_2-q_1q_2-q_1^2)-30q_2"

"\u03c0(q_1,q_2)=90q_2-q_2q_1-q_2^2" (7)

Derivative of "\u03c0" with respect to "q_1" for (6)

"d\u03c0\/(dq_1 )=90-2q_1-q_2=0"

"90-q_2=2q_1"

"q_1=(90-q_2)\/2" (8)


Derivative of "\u03c0" with respect to "q_2" for (7)

"d\u03c0\/(dq_1 )=90-q_1-2q_2-=0"

"90-q_1=2q_2"

"q_2=(90-q_1)\/2" (9)

Substitute (9) into (8)

"q_1=(90-(90-q_1)\/2)\/2"

"q_1=(90+q_1)\/4"

"90+q_1=4q_1"

"90=3q_1"

"q_1=30" units


Substitute "q_1" with 30 in (9)

"q_2=(90-30)\/2=60\/2"

"q_2=30" units


Each Firm Profit

Substitute the value of "q_1" and "q_1" into (6) and (7)

Firm 1 profit:

"\u03c0(q_1,q_2)=90(30)-(30)^2-30(30)=" $"900"

Firm 2 profit:

"\u03c0(q_1,q_2)=90(30)-30(30)-(30)^2=" $"900"